# M-LLM Based Video Frame Selection for Efficient Video Understanding

### v1

@@ 4.1 Experiment Setup
Training data We compile the training dataset from three sources: 1) 800K data from VideoChat2 [16], 2) 125K data from the TimeIT dataset [33], and 3) 178K data from the LLaVA-Video-178K dataset [56]. For the visual instruction tuning task, we use the entire training dataset. For the importance score prediction task, we use 400K video QA data where the video length exceeds 5 seconds.
Pseudo label generation we utilize Qwen2-VL-7B [41] to generate importance scores for each frame in the spatial pseudo labeling and concise captions for all frames in the temporal pseudo labeling. We use GPT-4o mini to propose the most helpful frames given the concise captions in the multi-frame evaluation.
Evaluation benchmarks Since our framework selects frames for video question answering, we evaluated its performance on benchmarks consisting of relatively longer videos, including open-ended video QA on ActivityNet-QA [51], and multi-choice QA on NExT-QA [47] and EgoSchema [27], LongVideoBench [46].
@@
Table 1: Comparison on open-ended question answering evaluation ActivityNet QA. Results with the “+ Selector” are ours.
@@ 4.2 Comparison with SOTA Video-LLMs
Table 2: Comparison on multi-choice question answering evaluation NExT-QA. Results with the “+ Selector” are ours.
@@
Table 3: Comparison on multi-choice question answering evaluation EgoSchema. Results with the “+ Selector” are ours.
Table 1, Table 2 and Table 3 present the comparison on ActivityNet-QA [51], NExT-QA [47] and EgoSchema [27] respectively. Following existing practice [26], performance on ActivityNet-QA is measured as “accuracy/correctness” metrics, and both metrics are evaluated using GPT-3.5, with higher values indicating better performance. Performances on NExT-QA and EgoSchema are the accuracy of of multi-choice questions where each question has 5 options. We prefill the word “Option” as the initial token to generate, and then use the next generated token as the prediction result.
@@ 4.3 Ablation Studies
We perform ablation studies on several components of the frame selection framework. First, we analyze the performance of the following frame selection methods on the video QA task:
  • Uniform sampling: the default uniform sampling
@@
We further show that the video QA system can use fewer frames to reason videos with the help of our frame selector. Table 5 shows the performance of LLaVA-NeXT-Video on NExT-QA taking different number of frames as input and the inference speed of LLaVA-NeXT-Video with different number of input frames. The inference speed was measured with float16 precision using a batch size of 1 on a single A100 GPU, employing the Hugging Face implementation. When taking the same number of frames as input, our framework incurs additional inference costs to select frames using the M-LLM selector. However, this increase in inference time is not significant thanks to the efficient design of the frame selector. Moreover, the frames selected by the M-LLM selector are more useful for answering the question. Therefore, the model using a selector with ${n}$-frame input can achieve similar video QA performance as the model without a selector using ${2n}$-frame input. For example, the configure $128\rightarrow 4$ outperforms 8-frame uniform sampling with a faster inference speed.
@@
As discussed in Section 3.2, one advantage of our framework is that the frame selector features a lightweight design. We examines two key hyperparameters that influence the computational efficiency: the number of tokens to represent a frame and the size of the base LLM. By default, we use Qwen2.5 1.5b as the LLM backbone and use 9 tokens for video frame representation. Table 6 and Tabel 7 present the ablation studies examining these two factors. “no selector” indicates sample the video frames uniformly.
@@ 4.4 Visualization of selected frames
Figure 4 presents two examples of selected frames from the video conditioned the question. Each question involves two events and thus requires temporal reasoning. Estimating the importance of a single frame is challenging without reference to prior or subsequent frames. The frame selector effectively identifies frames containing the answers to the questions.
@@ 5 Conclusion
In this paper, we propose a lightweight M-LLM-based frame selector to improve both performances and efficiency in video QA. This selector is question-aware and takes dense video frames and the question as input, selecting the most relevant frames for answering the question. We can then use any multi-image or video M-LLMs with the selected frames to complete the question answering. To train the frame selector, we introduce spatial and temporal pseudo-labeling due to the limited public annotations for video frame-level importance. Our experiments on two medium-length video QA benchmarks (ActivityNet QA and NExt-QA) and two long-video QA benchmarks (EgoSchema and LongVideoBench) demonstrate the effectiveness of our proposed method.
@@ Appendix A Prompt for Pseudo Label Generation
the video, each with a corresponding caption. Below is
the list of frames and their captions:

### v2

@@ 4.1 Experiment Setup
Training data We compile the training dataset from three sources: 1) 800K data from VideoChat2 [20], 2) 125K data from the TimeIT dataset [38], and 3) 178K data from the LLaVA-Video-178K dataset [63]. For the visual instruction tuning task, we use the entire training dataset. For the importance score prediction task, we use 400K video QA data where the video length exceeds 5 seconds. Pseudo label generation we utilize Qwen2-VL-7B [46] to generate importance scores for each frame in the spatial pseudo labeling and concise captions for all frames in the temporal pseudo labeling. We use GPT-4o mini to propose the most helpful frames given the concise captions in the multi-frame evaluation. Evaluation benchmarks Since our framework selects frames for video question answering, we evaluated its performance on benchmarks consisting of relatively longer videos, including open-ended video QA on ActivityNet-QA [58], and multi-choice QA on NExT-QA [54], VideoMME [9], EgoSchema [32], LongVideoBench [53].
@@
Table 1: Comparison of open-ended question answering evaluation on ActivityNet QA. Results with the “+ Selector” are ours.
@@ 4.2 Comparison with SOTA Video-LLMs
Table 2: Comparison of multi-choice question answering evaluation on NExT-QA. Results with the “+ Selector” are ours.
@@
Table 3: Comparison of multi-choice question answering evaluation on EgoSchema. Results with the “+ Selector” are ours.
Model short medium long average
Qwen2-VL (baseline) 69.1 53.0 51.6 58.1
Qwen2-VL + our selector 69.6 54.1 51.9 58.7 (0.6$\uparrow$)
Table 4: Performances on VideoMME. “+ Selector” are ours.
Table 1, Table 2, Table 3 and Table 4 present the comparison on ActivityNet-QA [58], NExT-QA [54] and EgoSchema [32], VideoMME [9] respectively. Following existing practice [31], performance on ActivityNet-QA is measured as “accuracy/correctness” metrics, and both metrics are evaluated using GPT-3.5, with higher values indicating better performance. Performances on NExT-QA and EgoSchema are the accuracy of of multi-choice questions where each question has 5 options. We prefill the word “Option” as the initial token to generate, and then use the next generated token as the prediction result.
@@ 4.3 Ablation Studies
We conduct ablation studies on several components of the frame selection framework. First, we analyze the performance of the following frame selection methods in the video QA tasks:
  • Uniform sampling: the default uniform sampling.
@@
Not using video-LLM due to worse pseudo-label quality. We refer to Table 3 in [8] for justification. In the temporal reasoning tasks, although video-LLMs are trained with frames sampled from the same video, they fall behind image-trained M-LLMs. It implies image-trained M-LLMs could lead to better temporal pseudo-labels for our method. We tried to use LLaVA-NeXT-Video to generate pseudo-labels. It is more time-consuming but worse pseudo-labels quality. On NExT-QA, using frames labeled by Qwen2-VL and LLaVA-NeXT-Video are 63.9% and 62.8% respectively.
@@
We further show that the video QA system can use fewer frames to reason videos with the help of our frame selector. Table 6 shows the performance of LLaVA-NeXT-Video on NExT-QA taking different number of frames as input and the inference speed of LLaVA-NeXT-Video with different number of input frames. The inference speed was measured with float16 precision using a batch size of 1 on a single A100 GPU, employing the Hugging Face implementation. When taking the same number of frames as input, our framework incurs additional inference costs to select frames using the M-LLM selector. However, this increase in inference time is not significant thanks to the efficient design of the frame selector. Moreover, the frames selected by the M-LLM selector are more useful for answering the question. Therefore, the model using a selector with ${n}$-frame input can achieve similar video QA performance as the model without a selector using ${2n}$-frame input. For example, the configuration $128\rightarrow 4$ outperforms 8-frame uniform sampling with a faster inference speed.
@@
As discussed in Section 3.2, one advantage of our framework is that the frame selector features a lightweight design. We examines two key hyperparameters that influence the computational efficiency: the number of tokens to represent a frame and the size of the base LLM. By default, we use Qwen2.5 1.5b as the LLM backbone and use 9 tokens for video frame representation. Table 7 and Tabel 8 present the ablation studies examining these two factors. “no selector” indicates sample the video frames uniformly.
@@ 4.4 Visualization of selected frames
Qualitative Results: Figure 4 presents two examples of selected frames from the video conditioned the question. Each question involves two events and thus requires temporal reasoning. Estimating the importance of a single frame is challenging without reference to prior or subsequent frames. The frame selector effectively identifies frames containing the answers to the questions. More results in Appendix Section C .
Evaluate on Video Grounding Benchmarks: We compare the moment retrieval performance with SeViLa on QVHighlights. Ours achieves 43.9% R1@5 and 32.3% R1@7 while SeViLa achieves 54.5% R1@5 and 36.5% R1@7. Although our frame selector has lower performance than SeViLa, it is not specifically trained on QVHighlights. In other words, the comparable performance proves the overall correctness of the frame selection.
@@ 5 Conclusion
In this paper, we propose a lightweight M-LLM-based frame selector to improve both performances and efficiency in video QA. This selector is question-aware and takes dense video frames and the question as input, selecting the most relevant frames. We can then use any multi-image or video M-LLMs with the selected frames to complete the question-answering. To train the frame selector, we introduce spatial and temporal pseudo-labeling due to the limited public annotations for video frame-level importance. Our experiments on two medium-length video QA benchmarks (ActivityNet QA and NExt-QA) and two long-video QA benchmarks (EgoSchema and LongVideoBench, VideoMME) demonstrate the effectiveness of our proposed method.
@@ Appendix A Prompt for Pseudo Label Generation
the video, each with a corresponding caption. The question
I need to answer is:
{question}
Below is the list of frames and their captions: